OpenStudy (anonymous):

Given the point (4, 5) and the slope of 6, find y when x = 24. a. 118 b. 144 c. 125 d. 149

OpenStudy (anonymous):

Using point-gradient formula: y-y₁ = m(x-x₁) y-5 = 6(x-4) y = 6(x-4) + 5 When x = 24: y = 6×(24-4) + 5 y = 6×20 + 5 y = 125, so C is correct.